probability

Assume that 13% of people are left-handed. If we select 10people at random, answer the following questions:

a. What is the probability that there are 3lefties in the group? correct check mark

b. What is the probability that there are atleast 3 lefties in the group? correct check mark

c. Find the expected number of right-handers inthe group. correct check mark

d. Find the standard deviation of the number ofright-handers in the group. correct check mark


e.
What is the probability that they are not allright-handed?
it would be 1 - P(0 left handers) ... howwould you find P(0 left handers)??
g. What is the probability that the majorityis right-handed?wouldn't this be P (x >= 6) = 1 - P(x<= 5) .. im gettin 1 - 0.999 which can't beright
Answer #1
e. P(not all right-handed) = P(at least 1 left-handed) = 1-P(0left-handed)
f. P(5 righties &5 lefties) = P(5 lefties) (this alreadyimplies there are 5 righties)
g. P(at least 6 righties)

answered by: Gabriel